Twice the difference of a number and two is six. Find the number.
I tried 2x-2=6
+2 +2
2x=8
2x/2 8/2

Answer by ankor@dixie-net.com(22740)

Twice the difference of a number and two is six.
Should be
2(x-2) = 6
divide both sides by 2
x - 2 = 3
x = 3 + 2
x = 5